OpenStudy (anonymous):

Rationalize the denominator and simplify 5 + square root 2 / 8 - square root 2

OpenStudy (mathmale):

\[\frac{ 5+\sqrt{2} }{ 8-\sqrt{2} }\]

OpenStudy (mathmale):

... is where we're starting from. To rationalize the den., please multiply both numerator and denominator of this fraction by the CONJUGATE of 8-Sqrt(2). That CONJUGATE is \[8+\sqrt{2}\]

OpenStudy (mathmale):

If necessary, draw your work, so that I can easily see what you've done.

OpenStudy (mathmale):